Status Report

NASA Space Station On-Orbit Status 18 Mar 2004

By SpaceRef Editor
March 18, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 18 Mar 2004

ISS systems continue to function nominally, except those noted previously or below. 

Update on Elektron:   The crew removed the Elektron’s Liquid Unit #6 (BZh-6), containing the O2 generator’s main subsystems, and replaced it with the remaining spare, BZh-7.  All parts contaminated by the electrolyte (alkaline solution: 25% potassium hydroxide [KOH] plus 75% water, toxicity level 2 [detrimental to health for long-term exposure]), were triple-bagged for disposal.  External surfaces were to be washed down thoroughly.  The crew wore latex gloves, mouth/noise masks and goggles.  Tomorrow, the crew will flush the Elektron with water and then attempt its activation.

Update on SFOG (TGK) candles:   As of tonight, a total of 11 candles have been burned since 3/13, all of them without failure.  There are 110 candles left on board, for ~55 days (assuming 100% functionality).  A new design is currently in testing, scheduled to enter service late this year (Progress 16P, possibly sooner).  According to TsUP/Moscow, one of the two onboard TGK combustion containers failed this morning, caused by its pump.  At this time, candles can be burned only sequentially, with a time gap of 3 hrs, not in parallel.  A spare pump is on board, and R&R will be scheduled ASAP.

Update on CMG-3:   A CMG anomaly investigation team met today at MCC-H and is expected to develop recommendations re the recent vibration signature of control moment gyro #3.  Early yesterday, CMG-3 vibrations showed a very brief spike exceeding the 0.06g limit, after a 0.05g episode earlier.  The team is looking into possible relationships to the LoCs (losses of communications) experienced over last weekend with CMG-2.

The PCS (portable computer system) laptop in the U.S. Airlock failed to start up after its weekly reboot, indicating a corrupted HDD (hard disk drive).  Two spare HDDs are ready configured; the failed drive will be replaced pending crewtime availability.   [The Flight Rule-required minimum of two PCSs remains available, with one in the Lab and one in the SM.]

Before starting the Elektron R&R, CDR/SO Michael Foale took air samples in the station in response to the odor detected by the crew recently from the machine, using the U.S. GSC (Grab Sample Container) and the new DST (Dual Sorbent Tube) equipment.   [Houston and Moscow will jointly investigate the odor-causing compound(s).]

Mike also conducted routine maintenance on CSA-CP #1003 (compound specific analyzer-combustion products).  Later, before the Elektron activity, the FE took air check readings with CSA-CPs #1003, #1009, and #1010.

CDR/SO Michael Foale activated the MSG (Microgravity Science Glovebox) and supported startup of another session of the PFMI (Pore Formation & Migration Investigation) experiment.  PFMI ops continue tomorrow.   [Communications between MSG and its laptop (MLC) were successfully re-established by the Science Officer yesterday.  Today, he installed the PFMI-02 sample ampoule into the thermal chamber.  PFMI studies detrimental porosity formation and mobility during controlled directional solidification processing in a micro-G environment, utilizing a transparent material (Succinonitrile {SCN}), so that direct video observation and recording of pore generation and mobility during controlled directional solidification can be made.  SCN is particularly well suited for the proposed investigation because it is transparent, solidifies in a manner analogous to most metals, and has a convenient melting point.  With toxicity level 0, its material properties are well known and it has been successfully used in previous microgravity experiments.]

FE Alex Kaleri continued troubleshooting the Russian/German GTS (Global Timing System), “wringing out” its cable network.   [Using a spectrum analyzer and the “FHS View” software application installed yesterday on Laptop TP1, Kaleri took measurements for calibrating cables and checking cable continuity to the SUBA (onboard equipment control system).  GTS electronics are installed in the SM, with two transmitters broadcasting a time signal at 400.1 MHz (1 watt) and 1428 MHz (0.5 watt) straight downward (nadir) for worldwide clock synchronization.  The system has been having problems for some time.] 

Sasha performed his regular maintenance on the BIO-5 Rasteniya-2 (“Plants-2”) greenhouse.  Replenishment of its water supply as required has now been added to the Russian task list.   [Rasteniya studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.  Regular maintenance involves monitoring of seedling growth, humidity measurements, watering to moisten the substrate if necessary, and photo/video recording.]

Mike Foale successfully completed another proficiency training run with the Canadian SSRMS (Space Station Remote Manipulator System).   [To cut down on activity time, the original plan of testing the “sticky” PDGF (power & data grapple fixture) was changed, and the CDR was given the Go to maneuver directly from the current SSRMS position to the ISS portside survey position.]

Later, the ground-commanded portside survey of the station’s exterior with the SSRMS cameras got underway at around 1:30pm EST, to last for 7 hrs.

Mike did the routine maintenance of the SM’s SOZh environment control & life support system, incl. the toilet equipment (ASU).

The crew’s interactive TV event with middle schoolers in Gainesville, FL, yesterday drew rave reviews by the ground.   [“Based on the cheering in the background at Howard Bishop Middle School, we think it was quite a success.”]   

TsUP/Moscow confirmed the hard failure of the BUPT automatic current converter control of SM battery #8.   [Its replacement with a spare is scheduled for 4/6-7.  Seven 800A storage batteries remain fully functional in the SM, five (of 6) in the FGB.]

Today’s CEO (Crew Earth Observations) targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were SE Asian smoke, Gulf of Siam (Dynamic event.  The basin of Thailand has been mentioned repeatedly in the last few weeks for major smoke palls.  These were drifting out to sea, towards the ISS track: shooting obliques left towards the Gulf of Siam.  Limb shots are best for visualization of these aerosol events), Internal waves, Vietnam (large afternoon sunglint disc left of track should have yielded views of internal waves along the entire coast of Vietnam and in the South China Sea), Madagascar floods (Dynamic event.  Floodwaters from Cyclone Gafilo have killed dozens of people and caused damage to towns and crops in the northern half of the island.  The cloud has finally cleared so that images of major rivers and deltas can be obtained.  ISS tracked along the west coast where the island’s major rivers empty into the sea), Betsiboka R Delta, Madagascar (400mm-lens.  This long term monitoring site has yielded very interesting results in terms of the growth of new islands and expansion of old ones), Dust front, Cent. Atlantic (Dynamic event.  NOAA’s research vessel, the Ronald. H. Brown, has reached the Canary Islands.  ISS/CEO pictures of a prior dust outblow event, taken when the ship was in the central Atlantic, are being forwarded to the crew, and one image will be posted on a NASA website.  Looking left and right for obliques of another dust event.  The southern margin of the dust-laden air mass may coincide with a major cloud band.  The northern margin may be visible ~4.5 minutes later on this ascending pass.  Images of dust margins and the Earth’s limb are best).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of today, 2:20pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is Off.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 28.1; ppO2 (mmHg) — 140.0; ppCO2 (mmHg) — 3.7;
  • SM Transfer Compartment:  Pressure (mmHg) — 743; temperature (deg C) — 20.3.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 25.0.
  • Node:  Pressure (mmHg) — 740.1; temperature (deg C) — 24.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 742.49; temperature (deg C) — 25.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 742.69; temperature (deg C) — 28.2; shell heater temp (deg C) — 26.4, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 23.9
  • PMA-2:  Shell heater temp (deg C) — 9.7

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (suntracking) and bias-angled 43 deg. for drag reduction (“sun slicer”)
  • SM batteries:  Battery #8 is off-line; all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #6 is off-line (capacity restoration mode, ROM); battery #5 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).

Propulsion System:

  • Total propellant load available: 3948 kg (8704 lb) as of 3/5/04   [SM(755) + FGB(2534) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist), until 3/28.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • *S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, powered on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 3:56am EST [= epoch]):

  • Mean altitude — 367.7 km
  • Apogee — 374.8km
  • Perigee — 360.5 km
  • Period — 91.90 min.
  • Inclination (to Equator) —  51.6288 deg
  • Eccentricity — 0.0010603
  • Solar Beta Angle — ~46 deg (magnitude decreasing)
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 110 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 30406

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.